2. Imaging Inc., a developer of radiology equipment, has stock outstanding as follows: 22,000 shares of cumulative preferred 4% stock, $140 par, and 73,000 shares of $5 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $82,500; second year, $173,900; third year, $211,900; fourth year, $232,700.

Compute the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".

1st Year 2nd Year 3rd Year 4th Year
Preferred stock (dividend per share) $ $ $ $
Common stock (dividend per share) $ $ $ $
Answer #1

Stock Outstanding:

Cumulative preferred 4% stock,$140 par = 22,000

Total preferred share capital=140*22000=$3,080,000

Common shares, $5 par = 73000shares

1st year:

Dividend distributed in first year=$82,500

Dividend due on preferred stock=4%*$3,080,000=$123,200

Entire $82500 is distributed to preferred stockholders

Dividend per share on preferred stock=$82500/22000=$3.75

Dividend on common stock=Nil

SECOND YEAR

Arrear of dividend on preferred share=123,200-82500=$40,700

This years dividend due=$123,200

Total dividend due on cumulative preferred shares=$123,200+$40,700=$163,900

Actual dividend paid during the year=$173,900

Dividend paid to preferred shareholders=$163,900

Dividend paid to common shareholders:

=$173,900-$163,900=$10000

Dividend per share paid to preferred shareholder=$163,900/22000=$7.45

Dividend per share paid to common shareholder=$10000/73000=$0.14

THIRD YEAR

Dividend due for preferred shareholders=$123,200

Dividend distributed =$211,900

Dividend paid to common shareholders($211,900-123,200)=$88,700

Dividend per share paid to preferred shareholder=123,200/22,000=$5.6

Dividend per share paid to common shareholder=$88,700/73000=$1.22

FOURTH YEAR

Dividend due for preferred shareholders=$123,200

Dividend distributed =$232,700

Dividend paid to common shareholders($232,700-123,200)=$109,500

Dividend per share paid to preferred shareholder=123,200/22000=$5.6

Dividend per share paid to common shareholder=$109500/73,000=$1.5

YEAR WISE DIVIDEND PER SHARE

Dividend per share

Preferred

Common

First year

$3.75

0

Second Year

$7.45

$0.14

Third Year

$5.6

$1.22

Fourth Year

$5.6

$1.5

ANY DOUBTS OR CORRECTIONS?

JUST LEAVE A COMMENT BELOW

THANK YOU